Give the usual ax^2+bx+c form of the quadratic function which has a= 1, and has two zeros z= -5 and z= 3
Found 3 solutions by edjones, Mathtut, solver91311:
Answer by edjones(8007)

x+5=0, x-3=0
(x+5)(x-3)=0
x^2+2x-15
.
Ed

Answer by Mathtut(3670)
Answer by solver91311(24713)



If or

Then or

And these are then the two factors of your quadratic trinomial:





John